OpenStudy (anonymous):

Medal, fan, testimonial for quickest answer. Please Help. Find all solutions in the interval [0, 2π). cos2x + 2 cos x + 1 = 0 x = 2π x = π x = pi divided by four., seven pi divided by four. x = pi divided by two., three pi divided by two.

OpenStudy (anonymous):

\[\cos^2x + 2 \cos x+1\]

OpenStudy (anonymous):

@haleyelizabeth2017 @gottennis121

OpenStudy (anonymous):

@Dani_Rose

OpenStudy (anonymous):

Please help me @Dani_Rose @leticiau @Preetha

OpenStudy (dani_rose):

idk how to do this but maybe this will help https://www.youtube.com/watch?v=CnhmIbagc04

OpenStudy (anonymous):

@Jaynator495 @KendrickLamar2014 @Preetha

OpenStudy (anonymous):

@jabez177 @imqwerty

OpenStudy (anonymous):

@Marie1738 @Qwertty123

OpenStudy (anonymous):

Please Help Me! These are my options: x = 2π x = π x = pi divided by four., seven pi divided by four. x = pi divided by two., three pi divided by two.

OpenStudy (anonymous):

Im not looking for answer, I just need help solving

OpenStudy (marie1738):

Okay. So I suck at math, but English. tag me anytime... Just some knowledge you should know beforehand..... BUT. I am very capable of doing research online, sooo.. Here are some explanations I found online in reference to this particural problem: #1: Basically the solutions of this particular equation is going to be restricted to one revolution. So first: (cosx+1)^2 = cos^2(x)+2cosx+1=0 cosx+1=0 cosx = -1 (cosines equal -1 when x=pi, and pi only occurs once from 0 to pi) x=pi #2: we need the trig identities to rescue us! cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 here the 2nd does the trick: subbing in, we get 2 cos^2(x) - 1 + 2cosx + 1=0 or 2 cos^2(x) + 2cosx=0 cos^2(x) + cosx = 0 cosx(cosx+1)=0 so cosx = 0 or cosx = -1 this happens at pi/2 and 3pi/2 #3: you can factor as follows: (cos x + 1)^2 = 0 cos x = -1 This occurs when x = pi.

OpenStudy (marie1738):

I do hope I helped. Even if I did not personally know the answer to the problem

OpenStudy (anonymous):

Thank you so much! Im not understanding much but I will look into it. I think it is saying that it happens at pi/2 and 3pi/2

OpenStudy (marie1738):

I can't tell you, but the answer is in fact B. x= pi

OpenStudy (marie1738):

Have a great day young friend

OpenStudy (anonymous):

Thanks! You too!